OpenStudy (anonymous):

Given the system of equations, what is the value of the system determinant? 2x + y = 8x - y = 10 a.)-1 b.)-2 c.)-3

OpenStudy (amistre64):

2x + y = 8 x - y = 10 | 2 1 | | 1 -1 |

OpenStudy (anonymous):

i'm not sure what a system determinent is?

OpenStudy (amistre64):

it is prolly best explained by your material

OpenStudy (amistre64):

it is the determinant of the coefficient matrix that correspond to your x and y valriables

OpenStudy (phi):

Your 2 equations are as amistre typed we can write this as a matrix equation: \[\left[\begin{matrix}2 & 1 \\ 1 & -1\end{matrix}\right]\left[\begin{matrix}x \\ y\end{matrix}\right]=\left[\begin{matrix}8 \\ 10\end{matrix}\right]\]

OpenStudy (phi):

They want you to find the determinant of the matrix, the "system determinant"